# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 10. Numerical experiment
where all the primes stand for the same derivative $\partial_{y_{l}}$, $1\leq l\leq n$, and $\nabla g=(\partial_{y_{1}}g,\dots,\partial_{y_{n}}g)^{T}$. The fraction on the right is $O(\epsilon^{s-1})$, therefore $\|g_{\epsilon}\|_{C^{1}(V)}<c$, $0<\epsilon\ll 1$.
To prove $\|g_{\epsilon}^{\prime}\|_{C^{s-1}(V)}<c$, we argue similarly to (B.8). The case $h\geq\epsilon$ is completely analogous. If $0<h<\epsilon$, from (B.11) and the exactness of $\varphi$:
@@
(B.12) $\begin{split}&\Delta g_{\epsilon}^{\prime}(y,h)-\Delta g^{\prime}(y,h)\\ &=\sum_{j}\bigg{[}\varphi^{\prime}\bigg{(}\frac{y_{h}}{\epsilon}-j\bigg{)}-\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\bigg{]}\frac{g(\epsilon j)-[g(y_{h})+(\epsilon j-y_{h})\nabla g(y_{h})]}{\epsilon}\\ &\quad+\sum_{j}\varphi^{\prime}\bigg{(}\frac{y}{\epsilon}-j\bigg{)}\frac{\epsilon j-y}{\epsilon}[\nabla g(y_{h})-\nabla g(y)].\end{split}$
@@
Dividing the first sum by $h^{s-1}=(h/\epsilon)^{s-1}\epsilon^{s-1}$ we get similarly to (B.10) that the ratio is bounded. The second sum divided by $h^{s-1}$ is bounded, because $g^{\prime}\in C^{s-1}(V)$. The proof is complete.
@@
The idea of the proof is to reduce $f_{\epsilon}$ to $G_{\epsilon}$ by first neglecting smooth terms in $f_{\epsilon}$ and then by neglecting lower order terms of finite smoothness, which still do not contribute to the DTB. This is done in sections C.1 and C.2, respectively.
@@ C.1. Simplification of $f_{\epsilon}$ by subtracting smooth terms
Let $\mathcal{B}_{\epsilon}\in L^{-3}(\mathcal{U}_{b})$ be a local parametrix for $\mathcal{T}_{\epsilon}$ on the open set $\{x\in\mathcal{U}_{b}:\chi(x)>1/2\}\supset\mathcal{U}_{0}$, see Figure 4. Below, we omit the part “$x\in\mathcal{U}_{b}$” in similar sets. Then
(C.2) $\begin{split}&f_{\epsilon}(x)=(\mathcal{B}_{\epsilon}F_{\epsilon})(x)+(\mathcal{H}_{\epsilon}f_{\epsilon})(x),\\ &\tilde{B}_{\epsilon}(x,\xi)-\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}\in S^{-1}(\mathcal{U}_{0}),\end{split}$
where $\tilde{B}_{\epsilon}\in S^{-3}(\mathcal{U}_{b})$ is the complete symbol of $\mathcal{B}_{\epsilon}$, and $\mathcal{H}_{\epsilon}:\mathcal{E}^{\prime}(\mathcal{U}_{b})\to C^{\infty}(\mathcal{U}_{0})$ is an operator with a smooth kernel, $H_{\epsilon}(x,z)$.
Figure 4. Illustration of the sets $\mathcal{U}_{0}$, $\mathcal{U}_{b}$, $\mathcal{U}$, $\{\chi(x)=1\}$, and the level set $\{\chi(x)=c\}$, $10<c<$ 1.
To prove the latter claim, use a partition of unity to write $f=f_{1}+f_{2}$, where
$\text{supp}\,f_{1}\subset\{\chi(x)>1/2\},\quad\text{supp}\,f_{2}\subset\{\chi(x)<3/4\}.$
Then $\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}(f_{1}+f_{2})=f_{1}+\mathcal{H}_{\epsilon}^{\prime}f_{1}+\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}f_{2}$, where $\mathcal{H}_{\epsilon}^{\prime}\in L^{-\infty}(\{\chi(x)>1/2\})$. The first two terms on the right follow by the properties of a parametrix. The fact that
$\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}:\mathcal{E}^{\prime}(\{\chi(x)<3/4\})\to C^{\infty}(\mathcal{U}_{0})$
@@
Derivatives of $H_{\epsilon}(x,z)$ with respect to $x$ and $z$ are uniformly bounded with respect to $x\in\mathcal{U}_{0}$, $z\in\mathcal{U}_{b}$, and $0<\epsilon\ll 1$. This follows from the way a symbol of $\mathcal{B}_{\epsilon}$ is constructed (see [52, Ch. I, eqs. (4.39), (4.40)]) and observing that all the seminorms of $\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}$ as member of $S^{0}(\mathcal{U}_{b})$ (i.e., the minimal constants $c_{m}$ and $c_{m_{1},m_{2}}$ in (2.7), where $r=0$) are uniformly bounded for $0<\epsilon\ll 1$.
@@ C.2. Extracting the leading local singularity of $f_{\epsilon}$
In the preceding section we converted $f_{\epsilon}$ to $\mathcal{B}_{\epsilon}F_{\epsilon}$ modulo a smooth function. We now represent $\mathcal{B}_{\epsilon}=\mathcal{B}_{\epsilon}^{(0)}+\Delta\mathcal{B}_{\epsilon}$ with some convenient $\mathcal{B}_{\epsilon}^{(0)}\in L^{0}(\mathcal{U}_{b})$, $\Delta\mathcal{B}_{\epsilon}\in L^{-1}(\mathcal{U}_{b})$. Then we use the mapping properties of $\Delta\mathcal{B}_{\epsilon}$ to show that $\Delta\mathcal{B}_{\epsilon}F_{\epsilon}$ does not contribute to the DTB.
Using the results in [53, Section 6.2, Chapter VIII] and [11, Proposition 4.2.4, Section 4.2] and (3.5), the principal symbol of $\mathcal{R}^{*}\mathcal{R}$ is given by
(C.4) $\begin{split}\frac{2\pi}{|\nu|}\frac{W^{2}(x,y)}{|\Delta_{\Phi}(x,\alpha)|},\ y=y(x,\xi),\nu=\nu(x,\xi),\end{split}$
where the functions $y(x,\xi)\in\mathcal{V}$ and $\nu(x,\xi)\in\mathbb{R}$ are obtained by solving the equations:
(C.5) $p=\Phi(x,\alpha),\ \xi=-\nu\text{d}_{x}\Phi(x,\alpha).$
The existence of a solution follows from assumption 3.2(3), and its local uniqueness follows from assumption 3.2(1).
@@
Thus we can find $\Delta\tilde{B}_{\epsilon}\in S^{-1}(\mathcal{U}_{b})$ (with all seminorms uniformly bounded when $0<\epsilon\ll 1$) such that
@@
Note that (LABEL:tilde_B) is required to hold only for $x\in\mathcal{U}_{0}$. Since $\mathcal{B}_{\epsilon}$ is a parametrix for $T_{\epsilon}$ on $\{\chi(x)>1/2\}$, we can select $\tilde{B}_{\epsilon}$ in (C.2) and $\Delta\tilde{B}_{\epsilon}$ so that
(C.10) $\begin{split}\tilde{B}_{\epsilon}(x,\xi)\equiv 0,\ \Delta\tilde{B}_{\epsilon}(x,\xi)\equiv 0,\ x\in\{\chi(x)<1/4\}.\end{split}$
@@
(C.11) $\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}F_{\epsilon}=\mathcal{W}_{\epsilon}g_{\epsilon},\ \mathcal{W}_{\epsilon}:=\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}\text{Op}(\chi(x)|\xi|)\chi(x)\mathcal{R}^{*}.$
By construction, $\mathcal{W}_{\epsilon}$ is an FIO of order $-1/2$, $\mathcal{W}_{\epsilon}\in I^{-1/2}(\mathcal{U}\times\mathcal{V}\times\mathcal{C}^{*})$. The canonical relation $\mathcal{C}^{*}$ is the image of $\mathcal{C}$ (see (3.3)) under the exchange of $T^{*}\mathcal{U}$ and $T^{*}\mathcal{V}$. By (LABEL:tilde_B), all the seminorms of a symbol of $\mathcal{W}_{\epsilon}$, $W_{\epsilon}\in S^{0}(\mathcal{U}\times\mathcal{V}\times\mathbb{R})$, $0<\epsilon\ll 1$, are bounded.
Recall that $B_{\infty,\infty}^{s}(\mathbb{R}^{n})=C^{s}(\mathbb{R}^{n})$, $s>0$, $s\not\in\mathbb{N}$ [1, Remark 6.4(2)]. We use [23, Proposition 5.5] with $p,q=\infty$, which asserts that
(C.12) $\mathcal{W}_{\epsilon}:C^{s-(1/2)-m_{c}(\infty)}(\mathbb{R}^{n})\to C^{s}(\mathbb{R}^{n}),\ m_{c}(p)=-(n-1)\big{|}(1/p)-(1/2)\big{|},$
is continuous. Here $s-(1/2)-m_{c}(\infty)>0$ and $s>0$. See also [45, Theorem, p. 173]. For the validity of (C.12) it is required that the phase function of $\mathcal{W}_{\epsilon}$ be non-degenerate (see the paragraph following (3.2)) and the amplitude be compactly supported in $x$ (see (C.10)).
Clearly, $m_{c}(\infty)=-1/2$. Using (C.12) with $s=1/2$ and that $g_{\epsilon}\in C_{0}^{1/2}(\mathcal{V})$, $\|g_{\epsilon}\|_{C^{1/2}(\mathcal{U}_{b})}<c$, $0<\epsilon\ll 1$ (see Lemma 4.2), we conclude that $\mathcal{W}_{\epsilon}g_{\epsilon}\in C_{0}^{1/2}(\mathcal{U}_{b})$ and $\|\mathcal{W}_{\epsilon}g_{\epsilon}\|_{C^{1/2}(\mathcal{U}_{b})}<c$, $0<\epsilon\ll 1$. Therefore, by (2.3),
(C.13) $\|\mathcal{W}_{\epsilon}g_{\epsilon}\|_{L^{\infty}(\mathcal{U}_{b})}<c,0<\epsilon\ll 1;\ |\mathcal{W}_{\epsilon}g_{\epsilon}(x_{0}+\epsilon\check{x})-\mathcal{W}_{\epsilon}g_{\epsilon}(x_{0})|=O(\epsilon^{1/2})$
when $\check{x}$ is confined to any bounded set.
Combining (C.3) and (C.13) finishes the proof.
@@
(D.1) $H(\check{t},\check{q};\cdot)\equiv 0,\ \check{t}<-c.$
Also, for some smooth and bounded $\varkappa(t,q;\cdot)$,
(D.2) $\begin{split}&\partial_{\check{t}}^{k}\big{[}H(\check{t},\check{q};\cdot)-\varkappa(\epsilon(\check{t}+\check{q});\cdot)\check{t}^{-1/2}\big{]}=O(\check{t}^{-3/2}),\\ &k=0,1,\ \check{t}\to+\infty,\ |\check{t}|<A/\epsilon,\end{split}$
@@
(D.3) $\begin{split}&\partial_{\check{t}}\big{[}\varkappa(\epsilon(\check{t}+\check{q});\cdot)\check{t}^{-(1/2)}\big{]}=O(\check{t}^{-3/2}),\ \check{t}\to+\infty,\ |\check{t}|<A/\epsilon.\end{split}$
The big-$O$ terms are uniform with respect to $\check{q}$ and the second group of variables collectively denoted $\cdot$’, namely $\theta,\epsilon$. These variables are restricted to the sets indicated in Lemma 6.2.
For simplicity, we drop the arguments $\check{q}$, $\theta$ and $\epsilon$ from all the functions. The limit as $\check{t}\to-\infty$ in (6.16) follows immediately from (D.1). Next, suppose $\check{t}\to\infty$. Then
@@
where $\varkappa(\cdot)$ is the same as in (LABEL:H_props_2_a), (LABEL:H_props_2_b). In view of (D.1), the first of the integrals is split into five
(D.5) $\begin{split}&\bigg{(}\int_{-c}^{0}+\int_{0}^{1}+\int_{1}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{H_{1}(s)}{s-\check{t}}\text{d}s\\ &=J_{1}+\dots+J_{5}.\end{split}$
By (LABEL:H_props_2_a) with $k=0$, $H_{1}(s)=O(s^{-3/2})$, $s\to\infty$. It is trivial to see that $J_{1},J_{2}=O(1/\check{t})$. Breaking up the integral over $[1,\check{t}-1]$ into the integrals over $[1,\check{t}/2]$ and $[\check{t}/2,\check{t}-1]$ we get $J_{3}=O(1/\check{t})$. Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{5}=O(1/\check{t})$. Finally, using (LABEL:H_props_2_a) with $k=1$ and that $|s-\check{t}|\leq 1$, $\epsilon\check{t}\leq c$, we get
@@
which implies $J_{4}=O(\check{t}^{-3/2})$. Combining the five estimates proves that the first integral in (D.4) is $O(1/\check{t})$.
Consider the second integral in (D.4). Recall that $\check{t}\to\infty$. Write the integral in the form
(D.7) $\begin{split}J(\check{t},\epsilon):=&\int_{0}^{\infty}\frac{\chi_{1\text{D}}(\epsilon(\check{s}-\check{t}))\varkappa(\epsilon(\check{s}+\check{q}))}{(\check{s}-\check{t})\check{s}^{1/2}}\text{d}\check{s}=\epsilon^{1/2}\int_{0}^{\infty}\frac{\chi_{1\text{D}}(s-t)\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ =&\epsilon^{1/2}\int_{0}^{\infty}\frac{(\chi_{1\text{D}}(s-t)-\chi_{1\text{D}}(0))\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ &+\epsilon^{1/2}\chi_{1\text{D}}(0)\int_{0}^{\infty}\frac{\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s,\ q=\epsilon\check{q}.\end{split}$
The integral on the second line is clearly uniformly bounded. The identity [43, Eq. 2.2.5.26] with $\alpha=1/2$:
@@
Here the term $O(1)$ is uniform with respect to $t,q\in\mathbb{R}$. Therefore $J(\check{t},\epsilon)=O(\epsilon^{1/2})$. Combining the results proves (6.16).
@@
Assertion (D.1) is obvious from (6.11). To prove (LABEL:H_props_2_a), (LABEL:H_props_2_b), begin by proving the following auxiliary lemma.
@@
(D.10) $\begin{split}\partial_{t}^{k}\big{[}&\textstyle\sum_{j}\varphi(t+b-j)f(\epsilon j)(j-b)_{+}^{1/2}-f(\epsilon(t+b))t^{1/2}\big{]}=O\big{(}t^{-3/2}\big{)},\\ &t\to+\infty,\ \epsilon\to 0,\ \epsilon|t|<A,\ k=0,1,2.\end{split}$
The big-$O$ term is uniform in $b\in\mathbb{R}$.
@@
Denote $g(t):=f(\epsilon t)(t-b)^{1/2}$ and $u:=t+b$. Taylor expanding $g$ gives
(D.11) $\begin{split}&g(j)=g(u)+g^{\prime}(u)(j-u)+R_{1}(j,u),\\ &R_{1}(j,u)=\int_{u}^{j}g^{\prime\prime}(s)(j-s)\text{d}s.\end{split}$

### v2

@@ 10. Numerical experiment
where all the primes stand for the same derivative $\partial_{y_{l}}$ for some $l=1,2,\dots n$. The fraction on the right is $O(\epsilon^{s-1})$, therefore $\|g_{\epsilon}\|_{C^{1}(V)}<c$, $0<\epsilon\ll 1$.
To prove $\|g_{\epsilon}^{\prime}\|_{C^{s-1}(V)}<c$, we argue similarly to the case $s<1$. As before, if $|h|\geq\epsilon$, we only need to show that $|g_{\epsilon}^{\prime}(y)-g^{\prime}(y)|\leq c\epsilon^{s-1}$. The latter inequality immediately follows from (B.12).
If $|h|<\epsilon$, from (B.12) and the exactness of $\varphi$:
@@
(B.13) $\begin{split}&g_{\epsilon}^{\prime}(y_{h})-g_{\epsilon}^{\prime}(y)\\ &=\sum_{j}\bigg{[}\varphi^{\prime}\bigg{(}\frac{y_{h}-\epsilon j}{\epsilon}\bigg{)}-\varphi^{\prime}\bigg{(}\frac{y-\epsilon j}{\epsilon}\bigg{)}\bigg{]}\frac{g(\epsilon j)-[g(y)+\text{d}_{y}g(y)\cdot(\epsilon j-y)]}{\epsilon}.\end{split}$
@@
Dividing by $|h|^{s-1}=(|h|/\epsilon)^{s-1}\epsilon^{s-1}$ we get similarly to (B.11) that the ratio is bounded. The proof is complete.
@@
The idea of the proof is to reduce $f_{\epsilon}$ to $\mathcal{G}_{\epsilon}$ by first neglecting smooth terms in $f_{\epsilon}$ and then by neglecting lower order terms of finite smoothness, which still do not contribute to the DTB. This is done in Sections C.1 and C.2, respectively.
@@ C.1. Simplification of $f_{\epsilon}$ by subtracting smooth terms
Recall that $\chi\in C_{0}^{\infty}(\mathcal{U}_{b})$. Let $\mathcal{B}_{\epsilon}\in L^{-3}(\mathcal{U}_{b})$ be a local parametrix for $\mathcal{T}_{\epsilon}$ on the open set $\mathcal{U}_{1/2}:=\{x\in\mathcal{U}_{b}:\chi(x)>1/2\}\Supset\mathcal{U}_{0}$, see Figure 4. Below, we omit the part “$x\in\mathcal{U}_{b}$” in similar sets. Then
(C.2) $\begin{split}&f_{\epsilon}(x)=(\mathcal{B}_{\epsilon}F_{\epsilon})(x)+(\mathcal{H}_{\epsilon}f_{\epsilon})(x),\\ &\tilde{B}_{\epsilon}(x,\xi)-\chi_{1}(\xi)\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}\in S^{-1}(\mathcal{U}_{1/2}),\end{split}$
where
  1. (1)
    $\tilde{B}_{\epsilon}\in S^{-3}(\mathcal{U}_{b})$ is the complete symbol of $\mathcal{B}_{\epsilon}$, i.e. $\mathcal{B}_{\epsilon}=\text{Op}(\tilde{B}_{\epsilon})$,
  2. (2)
    $\mathcal{H}_{\epsilon}:\mathcal{E}^{\prime}(\mathcal{U}_{b})\to C^{\infty}(\mathcal{U}_{0})$ is an operator with a smooth kernel, $H_{\epsilon}(x,z)\in C^{\infty}(\mathcal{U}_{0}\times\mathcal{U}_{b})$, and
  3. (3)
    $\chi_{1}\in C^{\infty}(\mathbb{R}^{2})$ satisfies $\chi_{1}(\xi)\equiv 0$, $|\xi|\leq R$, and $\chi_{1}(\xi)\equiv 1$, $|\xi|\geq R+1$, for some $R\gg 1$.
Figure 4. The sets $\mathcal{U}_{0}$, $\mathcal{U}_{b}$, $\mathcal{U}$, $\{\chi(x)=1\}$, and the level set $\{\chi(x)=c\}$, $10<c<$ 1.
To prove (C.2), use a partition of unity to write $f_{\epsilon}=f_{\epsilon}^{(1)}+f_{\epsilon}^{(2)}$, where
(C.3) $\text{supp}\,f_{\epsilon}^{(1)}\subset\mathcal{U}_{1/2}=\{\chi(x)>1/2\},\quad\text{supp}\,f_{\epsilon}^{(2)}\subset\{\chi(x)<3/4\}.$
Then $\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}f_{\epsilon}=f_{\epsilon}^{(1)}+\mathcal{H}_{\epsilon}^{\prime}f_{\epsilon}^{(1)}+\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}f_{\epsilon}^{(2)}$, where $\mathcal{H}_{\epsilon}^{\prime}\in L^{-\infty}(\mathcal{U}_{1/2})$. The first two terms on the right follow by the properties of a parametrix, [55, Section I.4, Appendix]. By construction, $\text{dist}(\{\chi(x)<3/4\},\mathcal{U}_{0})>0$, and the property
(C.4) $\mathcal{B}_{\epsilon}\mathcal{T}_{\epsilon}:\mathcal{E}^{\prime}(\{\chi(x)<3/4\})\to C^{\infty}(\mathcal{U}_{0})$
@@
The derivatives $\partial_{x}^{m_{1}}\partial_{z}^{m_{2}}H_{\epsilon}(x,z)$, $m_{1},m_{2}\in\mathbb{N}_{0}^{2}$, are uniformly bounded with respect to $x\in\mathcal{U}_{0}$, $z\in\mathcal{U}_{b}$, and $0<\epsilon\ll 1$. This follows from the above discussion, the way a symbol of $\mathcal{B}_{\epsilon}$ is constructed (see [55, Ch. I, eqs. (4.39), (4.40)]), and the observation that all the seminorms of $\chi_{1}(\xi)\big{(}|\xi|\tilde{D}(x,\xi)+\kappa|\epsilon\xi|^{3}\big{)}^{-1}$ as member of $S^{0}(\mathcal{U}_{1/2})$ (i.e., the minimal constants $c_{m}$ and $c_{m_{1},m_{2}}$ in (2.7), where $r=0$) are uniformly bounded for $0<\epsilon\ll 1$. Likewise, the essence of the second line in (C.2) is that all the seminorms of the left-hand side as member of $S^{-1}(\mathcal{U}_{1/2})$ are uniformly bounded for $0<\epsilon\ll 1$ as well.
@@ C.2. Extracting the leading local singularity of $f_{\epsilon}$
In the preceding section we converted $f_{\epsilon}$ to $\mathcal{B}_{\epsilon}F_{\epsilon}$ modulo a $C^{\infty}(\mathcal{U}_{0})$ function. We now represent $\mathcal{B}_{\epsilon}=\mathcal{B}_{\epsilon}^{(0)}+\Delta\mathcal{B}_{\epsilon}$ with some convenient $\mathcal{B}_{\epsilon}^{(0)}\in L^{0}(\mathcal{U}_{b})$, $\Delta\mathcal{B}_{\epsilon}\in L^{-1}(\mathcal{U}_{b})$. Then we use the mapping properties of $\Delta\mathcal{B}_{\epsilon}$ to show that $\Delta\mathcal{B}_{\epsilon}F_{\epsilon}$ does not contribute to the DTB.
Using the results in [56, Chapter VIII, Section 6.2] and (3.5), the principal symbol of $\mathcal{R}^{*}\mathcal{R}$ is given by
(C.6) $\begin{split}\frac{2\pi}{|\nu|}\frac{W^{2}(x,y)}{|\Delta_{\Phi}(x,\alpha)|},\ y=y(x,\xi),\nu=\nu(x,\xi),\ (x,\xi)\in\mathcal{U}\times(\mathbb{R}^{2}\setminus 0_{2}),\end{split}$
where the functions $y(x,\xi)\in V$ and $\nu(x,\xi)\in\mathbb{R}$ are obtained by solving the equations:
(C.7) $p=\Phi(x,\alpha),\ \xi=-\nu\text{d}_{x}\Phi(x,\alpha),\ (\alpha,p)\in V.$
The existence of a local solution follows from Assumption 3.2(3). Its local uniqueness and smoothness follows from Assumption 3.2(1). In Theorem 3.6 we assume that the GRT data are not redundant, so only one solution exists. Note that the symbol in (C.6) is integrable at the origin $\xi=0$. More precisely, it satisfies the first inequality in (2.7) with $n=2$ and $\delta=1$.
@@
Thus we can find $\Delta\tilde{B}_{\epsilon}\in S^{-1}(\mathcal{U}_{b})$ (with all the seminorms uniformly bounded for $0<\epsilon\ll 1$) such that (see (C.2))
@@
(C.12) $\begin{split}&\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}F_{\epsilon}=\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon},\\ &\mathcal{B}_{1}:=\text{Op}\big{(}\Delta\tilde{B}_{\epsilon}(x,\xi)\big{)}\text{Op}(\chi(x)|\xi|)\chi(z)\in L^{0}(\mathcal{U}).\end{split}$
Arguing similarly to Section 6.1 (from the beginning through (6.7)), we get
(C.13) $\begin{split}&\|\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon}\|_{L^{\infty}(\mathcal{U}_{b})}<c,0<\epsilon\ll 1;\\ &|(\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon})(x_{0}+\epsilon\check{x})-(\mathcal{B}_{1}\mathcal{R}^{*}g_{\epsilon})(x_{0})|=O(\epsilon^{1/2}),\ \check{x}\in\mathcal{K},\end{split}$
where $\mathcal{K}$ is any compact set.
Combining (5.7), (C.2), (C.5), (LABEL:tilde_B) and (C.13) finishes the proof.
@@
(D.1) $H(\check{t},\check{q};*)\equiv 0,\ \check{t}<-c.$
Also, for some smooth and bounded $\varkappa(t,q;\theta)$,
(D.2) $\begin{split}&\partial_{\check{t}}^{k}\big{[}H(\check{t},\check{q};*)-\varkappa(\epsilon\check{t},\epsilon\check{q};\theta)\check{t}^{-1/2}\big{]}=O(\check{t}^{-3/2}),\\ &k=0,1,\ \check{t}\to+\infty,\ \check{t}<A/\epsilon,\end{split}$
@@
(D.3) $\begin{split}&\partial_{\check{t}}\big{[}\varkappa(\epsilon\check{t},\epsilon\check{q};\theta)\check{t}^{-1/2}\big{]}=O(\check{t}^{-3/2}),\ \check{t}\to+\infty,\ \check{t}<A/\epsilon.\end{split}$
The big-$O$ terms are uniform with respect to $\check{q}$, $\theta$, and $\epsilon$. These variables are restricted to the sets $|\check{q}|\leq A/\epsilon$, $\theta\in I_{\alpha}$, and $0<\epsilon\ll 1$.
For simplicity, we drop the arguments $\check{q}$, $\theta$ and $*=(x,\theta,\epsilon)$ from all the functions. From (D.1) and (LABEL:H_props_2_a) with $k=0$ we get $H(\check{t})=O(\check{t}^{-1/2})$, $s\to\infty$. From Assumption 3.4(1), (6.13), and (6.14), it follows that $H(\check{t},\check{q};*)$ and $\partial_{\check{t}}H(\check{t},\check{q};*)$ are uniformly bounded. The limit as $\check{t}\to-\infty$ in (6.19) easily follows from (6.17).
Next, suppose $\check{t}\to\infty$. Then
@@
where $\varkappa(\cdot)$ is the same as in (LABEL:H_props_2_a), (LABEL:H_props_2_b). Using (D.1), split the first of the integrals into four
(D.5) $\begin{split}&\bigg{(}\int_{-c}^{t/2}+\int_{t/2}^{\check{t}-1}+\int_{\check{t}-1}^{\check{t}+1}+\int_{\check{t}+1}^{\infty}\bigg{)}\chi_{1\text{D}}(\epsilon(s-\check{t}))\frac{H_{1}(s)}{s-\check{t}}\text{d}s\\ &=J_{1}+\dots+J_{4}.\end{split}$
By (LABEL:H_props_2_a) with $k=0$, $H_{1}(s)=O(s^{-3/2})$, $s\to\infty$. It is trivial to see that $J_{1}=O(1/\check{t})$ and $J_{2}=O(\check{t}^{-3/2}\ln\check{t})$. In the first integral we use $s-\check{t}\asymp\check{t}$, and in the second one we use that $s\asymp\check{t}$ to bound above $H_{1}(s)$ by $c\check{t}^{-3/2}$. Breaking up the integral over $[\check{t}+1,\infty)$ into the integrals over $[\check{t}+1,2\check{t}]$ and $[2\check{t},\infty)$ gives $J_{4}=O(\check{t}^{-3/2}\ln\check{t})$. Finally, using (LABEL:H_props_2_a) with $k=1$ and that $|s-\check{t}|\leq 1$, $\epsilon\check{t}\leq c$, we get
@@
which implies $J_{3}=O(\check{t}^{-3/2})$. Combining the four estimates proves that the first integral in (D.4) is $O(1/\check{t})$.
Consider the second integral in (D.4), which we denote $J(\check{t},\epsilon)$. Recall that $\check{t}\to\infty$. Write $J(\check{t},\epsilon)$ in the form
(D.7) $\begin{split}J(\check{t},\epsilon):=&\epsilon^{1/2}\int_{0}^{\infty}\frac{\chi_{1\text{D}}(s-t)\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ =&\epsilon^{1/2}\int_{0}^{\infty}\frac{(\chi_{1\text{D}}(s-t)-\chi_{1\text{D}}(0))\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s\\ &+\epsilon^{1/2}\chi_{1\text{D}}(0)\int_{0}^{\infty}\frac{\varkappa(s+q)}{(s-t)s^{1/2}}\text{d}s,\ q=\epsilon\check{q},\ t=\epsilon\check{t}.\end{split}$
It is not hard to show that the integral on the second line is uniformly bounded. Recall that $\chi_{1\text{D}}\in C_{0}^{\infty}(\mathbb{R})$. The identity [47, Eq. 2.2.5.26] with $\alpha=1/2$:
@@
Here the term $O(1)$ is uniform with respect to $0<t<A,|q|<A$. Therefore $J(\check{t},\epsilon)=O(\epsilon^{1/2})$.
Given that $\partial_{\check{t}}H$ is bounded, it is clear that $G$ is bounded with $\check{t}$ in compact sets. Combining all the results proves (6.19).
@@
Assertion (D.1) is obvious from (6.14) because $h$ is compactly supported. To prove (LABEL:H_props_2_a), (LABEL:H_props_2_b), begin by proving the following auxiliary lemma.
@@
(D.10) $\begin{split}\partial_{\check{t}}^{k}\big{[}&\textstyle\sum_{j}\varphi(\check{t}+\check{q}-j)f(\epsilon j)(j-\check{q})_{+}^{1/2}-f(\epsilon(\check{t}+\check{q}))\check{t}^{1/2}\big{]}=O\big{(}\check{t}^{-3/2}\big{)},\\ &\check{t}\to+\infty,\ \epsilon\to 0,\ \epsilon\check{t}<A,\ k=0,1,2.\end{split}$
The big-$O$ term is uniform in $|\check{q}|\leq A/\epsilon$.
@@
Denote $g(\check{t}):=f(\epsilon\check{t})(\check{t}-\check{q})^{1/2}$ and $u:=\check{t}+\check{q}$. The difference in brackets in (D.10) becomes $g_{\epsilon}(u)-g(u)$, where $g_{\epsilon}$ stands for the interpolated $g$ similarly to (3.14) (see (D.12)). Taylor expanding $g$ gives
(D.11) $\begin{split}&g(j)=g(u)+g^{\prime}(u)(j-u)+R_{1}(j,u),\ j-u\in\text{supp}\,\varphi,\\ &R_{1}(j,u)=\int_{u}^{j}g^{\prime\prime}(s)(j-s)\text{d}s.\end{split}$